Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, December 07, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 07, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/07/98 (Monday)

Segment 1 report for 12/7/98: Trenching continued on Lebec Road to EMP 21.55. On Tejon Bench, trench was excavated to EMP 18.07 and backfill was completed to 18.10. Backfill also took place from EMP 16.92 to 16.88. On Grapevine Rd., welding and x-ray took place from EMP 15.48 to 15.56. Shading occurred from EMP 16.35 to 16.25. Excavation of the Tejon Bench/I-5 bore began at EMP 16.49). Electrical work continued at Grapevine Pump Station. The bore operations continued at Stream Crossing #9 (EMP 15.35) and the California Aqueduct (EMP 12.28). At the Stream Crossing #4 bore (EMP 9.03), shoring was removed and backfill began. Pullboxes were installed on both sides of the Hwy. 166 bore (EMP 1.18) and backfill was taking place. At Emidio Pump Station, office equipment was being installed and calibration of the pumps continued.

Segment 2B (Castaic to Sylmar): Report for 12-7-98. Crews continued clean up of the directional drill site at Feedmill Rd. In addition, spoils were removed from the right of way and general clean up occurred at Weldon Cyn Rd. Welding crews completed the tie-in of the 20" pipe at Weldon Cyn Rd while coating and x-ray crews progressed south of the tie-in. Asphalt was completed to EMP 72.28, backfill to EMP 72.18 while fiber optic duct and shading were completed to near EMP 72.11. Fiber optic cable crews worked pulling cables between Sierra Hwy and Balboa Blvd.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 7 December 1998: 3A: The bore crew at Tuxford Street finished the removal of their equipment from the bore pit. Welders began work on the fiber optic conduit for this bore. Welding, x-ray, and coating crews continued work at Pacoima Wash Bridge, with welding complete on the south side of the bridge. Ditch crews continued excavating from the north side tie in of Pacoima. NIGHT work occurred at San Fernando and Hollywood Way Bridge crossings, focusing on chipping additional concrete out from the bridges. 3C: Backfill continued near Sepulveda Boulevard, as well as general clean up. Crews progressed with their work at the aqueduct crossing as well as the high speed channel cover. Final paving was done at Polk Street.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page